OpenStudy (anonymous):

I suck at this... Find the rate of change for this equation 5x + 3y = -2

OpenStudy (anonymous):

And I'd appreciate it if someone could explain this to me.

OpenStudy (anonymous):

dy/dx=-5/3 it,s true

OpenStudy (anonymous):

f(x)=5x+3y+2 f'(x)=5+3dy/dx so let f(x) =0 3dy/dx=-5 Sorry if i used wrong method.. i just want to help and check it>>>